two thirds of a number reduced by 11 is equal to 4 more than the number. Find the number.
Answer by kapilsinghi(68)

let the number = x
(2/3)*x - 11 = x + 4
2x/3 - 11 = x + 4
2x/3 = x + 15
2x = 3x + 45
x = -45
kapilsinghi123@gmail.com